Find the equation of the locus of the center of the circle which moves so that it is tangent to the y-axis and to the circle of radius one (1) with center at (2,0).


1
Expert's answer
2022-06-13T01:08:13-0400

Let "P(x,y)" be the center of a moving circle tangent to the y-axis and to a circle with a radius "1" with center at "(2,0)."

The distance from point "P" to the y-axis is "|x|". The distance from the point "P" to a center "(2,0)" is


"\sqrt{(x-2)^2+(y-0)^2}"

 The distance from the point "P" to a circle is 

"\sqrt{(x-2)^2+(y-0)^2}-1"


Then


"\sqrt{(x-2)^2+(y-0)^2}-1=|x|""\sqrt{(x-2)^2+(y-0)^2}=|x|+1""x^2-4x+4+y^2=x^2+2|x|+1""y^2=4x+2|x|-3"

Since "y^2\geq0, y\in \R," we take "|x|=x, x\geq0"

"y^2=4x+2x-3""y^2=6x-3"

The equation of the locus of the center of a moving circle tangent to the y-axis and to a circle with a radius 1 with center at "(2,0)" is the equaion of the parabola


"y^2=6x-3"